OpenStudy (pawanyadav):

Are the following function equal- (√x)^2. And. (x^2)^1/2

OpenStudy (pawanyadav):

I think not

hartnn (hartnn):

lets check, \(\sqrt x = x^{1/2}\) so 1st one is \(x^{(1/2) \times 2} = x \)

OpenStudy (pawanyadav):

Because I can't plug -1 in first But can plug in second function

hartnn (hartnn):

the second one is \(x^{(2 \times (1/2))} = x\) and both will always give out positive answers. so i don't see any difference

OpenStudy (pawanyadav):

This is not true for any negative integer

OpenStudy (pawanyadav):

Also the domain of both function are not same

OpenStudy (pawanyadav):

Which is a necessary condition for two function to be equal.

hartnn (hartnn):

yes, thats correct, since we are applying the square root function first, for the domain to be real, x should be >=0. you're right, tested with -4, the first one gives -4, the 2nd one will give 4

OpenStudy (pawanyadav):

Yea...

OpenStudy (pawanyadav):

The first function don't accept -4 or any negative number

ganeshie8 (ganeshie8):

\((\sqrt{-4})^2 = (2\sqrt{-1})^2 = (2i)^2 = 4i^2 = -4\)

ganeshie8 (ganeshie8):

\(((-4)^2)^{1/2} = 16^{1/2} = (4^2)^{1/2}=4\)

OpenStudy (pawanyadav):

But -4 is not in domain of √x But I think we don't considered I much in algebra

ganeshie8 (ganeshie8):

good point, so they don't even match at domain level

OpenStudy (pawanyadav):

iota ...I mean

hartnn (hartnn):

-4 is not in the real domain, if we can consider imaginary numbers, then -4 it is.